# A Customized SAT-based Solver for Graph Coloring

### v1

@@ 5 Conclusion
We presented ZykovColor, a new graph coloring algorithm using a SAT encoding that simulates the Zykov recurrence. We adopted the propagator presented in [19] and enhanced it by lazy addition reason clauses using the IPASIR-UP interface [12] of CaDiCal [4]. We also proposed using better lower bounds, a modified decision strategy of the underlying SAT solver CaDiCal [4].
In an ablation study we demonstrated the impact of the individual ideas. Most notably, we observed that using a propagator for pruning significantly improves the performance of the SAT solver. Additionally, the decision strategy has a significant impact on algorithm performance, with the strategy of CaDiCal performing better than a previous problem-specific strategy. Further, bottom-up optimization solves several more instances than top-down optimization, contrary to the results of gc-cdcl [19] where both configurations performed similarly.
Our experiments suggest that ZykovColor outperforms other graph coloring approaches from the literature on the DIMACS benchmark set. Experiments on the random Erdős-Rényi graph indicate that SAT-based methods still fall behind other methods like branch-and-price or DSatur on dense random graphs. This is a possible direction for further improving SAT-based methods. Lastly, ZykovColor could solve the yet unsolved DIMACS instance wap07 in 12 hours.
@@ Appendix A Detailed results on DIMACS instances
In this appendix we provide detailed results on the individual DIMACS instances.
@@
Table 2: Running time of algorithms for solved DIMACS instances. The best time for each instance is highlighted in bold.
Instance ZykovColor gc-cdcl POP-S Assignment CliColCom exactcolors Dsatur
1-FullIns_3 0.1 0.1 0.3 0.1 0.1 0.1 0.1
1-FullIns_4 0.1 0.1 0.4 0.1 0.1 - 0.1
1-FullIns_5 0.1 0.5 1.0 0.1 0.2 - 0.1
1-Insertions_4 175.7 - 1.2 0.4 1.5 - -
1-Insertions_5 - - - - - - -
1-Insertions_6 - - - - - - -
2-FullIns_3 0.1 0.1 0.2 0.1 0.1 0.1 0.1
2-FullIns_4 0.1 0.2 0.5 0.1 0.1 - 0.1
2-FullIns_5 1.1 186.3 4.1 0.2 2.3 - 1.6
2-Insertions_3 0.2 0.4 0.2 0.1 0.1 144.3 0.1
2-Insertions_4 - - - - - - -
2-Insertions_5 - - - - - - -
3-FullIns_3 0.1 0.1 0.2 0.1 0.1 0.1 0.1
3-FullIns_4 0.2 0.5 1.1 0.1 0.2 - 0.1
3-FullIns_5 5.9 14.8 23.2 5.2 32.2 - -
3-Insertions_3 0.9 7.0 0.2 0.1 0.1 - 2.3
3-Insertions_4 - - - - - - -
3-Insertions_5 - - - - - - -
4-FullIns_3 0.1 0.1 0.3 0.1 0.1 0.1 0.1
4-FullIns_4 0.2 1.9 2.3 0.1 0.8 - 0.1
4-FullIns_5 33.2 - 132.4 66.6 687.2 - -
4-Insertions_3 12.4 2300 0.3 0.1 0.2 - 2145
4-Insertions_4 - - - - - - -
5-FullIns_3 0.1 0.1 0.3 0.1 0.1 0.1 0.1
5-FullIns_4 0.2 3.8 5.9 0.2 5.9 - 0.3
abb313GPIA 350.2 - 24.9 0.7 - - -
anna 0.1 0.1 0.2 0.1 0.1 0.1 0.1
ash331GPIA 14.9 2.5 1.8 0.2 0.1 - 0.2
ash608GPIA 130.9 27.3 5.0 0.2 0.3 - -
ash958GPIA 1553 122.2 12.4 0.3 0.7 - 4.3
C2000.5 - - - - - - -
C2000.9 - - - - - - -
C4000.5 - - - - - - -
david 0.1 0.1 0.2 0.1 0.1 0.1 0.1
DSJC1000.1 - - - - - - -
DSJC1000.5 - - - - - - -
DSJC1000.9 - - - - - - -
DSJC125.1 0.6 0.8 0.3 0.1 0.1 - 0.1
DSJC125.5 - - - - - - -
DSJC125.9 - - - - - 7.9 -
DSJC250.1 - - - - - - -
DSJC250.5 - - - - - - -
DSJC250.9 - - - - - - -
DSJC500.1 - - - - - - -
DSJC500.5 - - - - - - -
DSJC500.9 - - - - - - -
DSJR500.1 0.1 0.1 1.5 0.1 0.1 305.4 0.1
DSJR500.1c 7.5 - - 3.9 36.9 637.5 -
DSJR500.5 78.4 - - 1044 31.0 3066 -
flat1000_50_0 - - - - - - -
flat1000_60_0 - - - - - - -
flat1000_76_0 - - - - - - -
flat300_20_0 - - - - - - -
flat300_26_0 - - - - - 1578 -
flat300_28_0 - - - - - - -
fpsol2.i.1 0.1 2.5 16.2 0.1 0.1 0.3 0.1
fpsol2.i.2 0.1 1.0 11.7 0.1 0.1 0.3 0.1
fpsol2.i.3 0.1 1.0 13.2 0.1 0.1 0.3 0.1
games120 0.1 0.1 0.3 0.1 0.1 0.1 0.1
homer 0.1 0.1 0.9 0.1 0.1 0.1 0.1
huck 0.1 0.1 0.3 0.1 0.1 0.1 0.1
inithx.i.1 0.2 3.5 21.4 0.2 0.2 1.0 0.1
inithx.i.2 0.2 2.0 18.7 0.2 0.1 0.3 0.1
inithx.i.3 0.2 1.9 19.3 0.1 0.1 0.3 0.1
jean 0.1 0.1 0.3 0.1 0.1 0.1 0.1
latin_square_10 - - - - - - -
le450_15a 24.1 13.2 4.1 0.3 0.2 - -
le450_15b 10.6 28.1 4.0 0.3 0.2 - -
le450_15c 647.8 - - - 58.7 - -
le450_15d 678.8 - - - 45.0 - -
le450_25a 0.2 0.5 4.9 0.2 0.1 2.1 0.1
le450_25b 0.2 0.3 4.9 0.2 0.1 1.8 0.1
le450_25c - - - - - - -
le450_25d - - - - - - -
le450_5a 5.9 30.3 2.0 0.1 0.1 - -
le450_5b 4.9 34.5 1.9 0.1 0.1 - -
le450_5c 2.6 5.6 3.0 0.1 0.1 - 0.1
le450_5d 2.1 5.1 3.0 0.1 0.1 - 11.2
miles1000 0.1 0.5 5.8 0.1 0.1 0.1 0.1
miles1500 0.1 1.9 21.9 0.1 0.1 0.1 0.1
miles250 0.1 0.1 0.3 0.1 0.1 0.1 0.1
miles500 0.1 0.1 0.4 0.1 0.1 0.1 0.1
miles750 0.1 0.2 2.3 0.1 0.1 0.1 0.1
mug100_1 0.1 0.1 0.3 0.1 0.1 1.2 -
mug100_25 0.1 0.1 0.3 0.1 0.1 1.1 -
mug88_1 0.1 0.1 0.3 0.1 0.1 0.7 371.7
mug88_25 0.1 0.1 0.3 0.1 0.1 0.7 241.7
mulsol.i.1 0.1 0.8 5.0 0.1 0.1 0.1 0.1
mulsol.i.2 0.1 0.4 5.4 0.1 0.1 0.1 0.1
mulsol.i.3 0.1 0.4 5.6 0.1 0.1 0.1 0.1
mulsol.i.4 0.1 0.4 5.7 0.1 0.1 0.1 0.1
mulsol.i.5 0.1 0.4 5.7 0.1 0.1 0.1 0.1
myciel3 0.1 0.1 0.3 0.1 0.1 0.1 0.1
myciel4 0.1 0.1 0.3 0.1 0.1 4.0 0.1
myciel5 0.1 0.1 0.7 0.1 0.5 - 1.2
myciel6 0.1 0.1 1981 0.1 932.6 - -
myciel7 0.2 0.5 - 0.2 - - -
qg.order100 - - - 1947 - - -
qg.order30 169.7 0.2 21.9 0.2 3.5 - 0.6
qg.order40 116.6 53.7 80.0 0.7 72.0 - -
qg.order60 1729 1710 2137 9.0 207.7 - -
queen10_10 555.8 - 371.3 - - 188.4 -
queen11_11 543.8 - 346.4 - - 1687 -
queen12_12 - - - - - - -
queen13_13 - - - - - - -
queen14_14 - - - - - - -
queen15_15 - - - - - - -
queen16_16 - - - - - - -
queen5_5 0.1 0.1 0.4 0.1 0.1 0.1 0.1
queen6_6 0.1 0.3 0.5 0.1 0.1 0.5 0.1
queen7_7 0.1 0.1 0.6 0.1 0.1 0.5 0.1
queen8_12 0.2 0.1 1.0 0.1 0.1 6.8 0.1
queen8_8 3.5 5.8 3.4 5.5 12.7 3.1 1.0
queen9_9 28.9 160.3 11.7 446.1 - 9.4 588.8
r1000.1 0.8 0.5 9.4 0.7 0.2 0.8 0.8
r1000.1c - - - - - - -
r1000.5 - - - 748.2 1870 - -
r125.1 0.1 0.1 0.3 0.1 0.1 0.1 0.1
r125.1c 0.1 3.5 37.9 0.1 0.1 0.1 0.1
r125.5 0.4 2.7 8.9 0.1 0.1 11.2 0.1
r250.1 0.1 0.1 0.5 0.1 0.1 0.1 0.1
r250.1c 0.2 40.3 103.3 0.1 0.3 28.9 1.0
r250.5 4.5 36.0 71.1 0.3 2.6 189.8 199.6
school1 0.8 3.1 18.2 0.9 0.6 935.0 5.7
school1_nsh 0.3 4.8 11.6 0.3 0.1 692.6 6.0
wap01a - - 1778 - 124.1 - -
wap02a - - 1054 - 144.4 - -
wap03a - - - - - - -
wap04a - - - - - - -
wap05a 0.9 5.6 61.2 0.9 0.7 5.9 0.9
wap06a 551.0 - 114.0 391.9 31.6 - -
wap07a - - - - - - -
wap08a - - 504.9 839.8 - - -
will199GPIA 0.4 1.2 3.2 0.2 0.2 3.8 0.3
zeroin.i.1 0.1 0.9 5.5 0.1 0.1 0.1 0.1
zeroin.i.2 0.1 0.4 4.3 0.1 0.1 0.1 0.1
zeroin.i.3 0.1 0.4 4.5 0.1 0.1 0.1 0.1
Table 2: Running time of algorithms for solved DIMACS instances. The best time for each instance is highlighted in bold.
Appendix B Memory usage on DIMACS and Erdős-Rényi Graphs
ZykovColor, Assignment, and gc-cdcl report their memory consumption, which we briefly discuss here. For the DIMACS graphs, all three algorithms used less than 4.5, 7.5, and 8.5 GB, respectively, except for qg.order100 (25, 25, and 44 GB respectively). On average, the algorithms use 385, 685, and 825 MB of memory, respectively. Comparing ZykovColor with Assignment, it is surprising that the latter encoding uses more memory than ZykovColor, despite the compactness of the assignment encoding. The Zykov-based encoding needs $O(n^{2})$ variables, one for each non-edge of the graph, which becomes quite large for large sparse graphs. Despite many such graphs in the benchmark set, ZykovColor requires noticeably less memory. This indicates a more effective traversal of the search tree and fewer conflict clauses being added so that the SAT solver requires less memory. For the small sizes of the Erdős-Rényi graphs, the algorithms ZykovColor, Assignment, and gc-cdcl all use about 60 MB of memory on average. However, one can observe that the memory of all three algorithms scales with the density of the graph. As the number of variables in Zykov-based encodings decreases with higher density, we would have expected the memory usage to decrease as well. But with the small graph sizes considered, this pattern might not be representative.

### v2

@@ 5 Conclusion
We presented ZykovColor, a new graph coloring algorithm using a SAT encoding that simulates the Zykov recurrence. We adopted the propagator presented in [19] and enhanced it with lazily added reason clauses using the IPASIR-UP interface [12] of CaDiCaL [4]. We also proposed a modified decision strategy of the underlying SAT solver CaDiCaL [4]. A decisive feature is the use of stronger lower bounds for pruning. For assignment or partial order-based SAT encodings, it would be unclear how to leverage them.
In an ablation study we demonstrated the impact of the individual ideas. The biggest improvement arises from using a propagator for pruning. Additionally, the decision strategy has a significant impact on algorithm performance. Here, the (modified) strategy of CaDiCaL performs better than the problem-specific strategy from [19]. Further, bottom-up optimization solves several more instances than top-down optimization, contrary to the results of gc-cdcl [19], where both configurations performed similarly.
ZykovColor outperforms other exact graph coloring approaches from the literature on the DIMACS benchmark set. On dense (random) Erdős-Rényi graphs, SAT-based methods still fall behind branch-and-price. This may motivate possible directions for further improving SAT-based methods.
@@ Appendix A Detailed Results on DIMACS Instances
In this appendix, we provide detailed results on the individual DIMACS instances (Table 2).
@@
Table 2: Instance data (number of vertices $n$, number of edges $m$ and density $d$) and running time of algorithms for solved DIMACS instances. The best time for each instance is highlighted in bold.
Instance $n$ $m$ $d$ ZykovColor gc-cdcl POP-S Assignment CliColCom exactcolors Dsatur
1-FullIns_3 30 100 0.23 0.1 0.1 0.3 0.1 0.1 0.1 0.1
1-FullIns_4 93 593 0.14 0.1 0.1 0.4 0.1 0.1 - 0.1
1-FullIns_5 282 3247 0.08 0.1 0.5 1.0 0.1 0.2 - 0.1
1-Insertions_4 67 232 0.10 158.1 - 1.2 0.4 1.5 - -
1-Insertions_5 202 1227 0.06 - - - - - - -
1-Insertions_6 607 6337 0.03 - - - - - - -
2-FullIns_3 52 201 0.15 0.1 0.1 0.2 0.1 0.1 0.1 0.1
2-FullIns_4 212 1621 0.07 0.1 0.2 0.5 0.1 0.1 - 0.1
2-FullIns_5 852 12201 0.03 0.8 186.3 4.1 0.2 2.3 - 1.6
2-Insertions_3 37 72 0.11 0.2 0.4 0.2 0.1 0.1 273.5 0.1
2-Insertions_4 149 541 0.05 - - - - - - -
2-Insertions_5 597 3936 0.02 - - - - - - -
3-FullIns_3 80 346 0.11 0.1 0.1 0.2 0.1 0.1 0.1 0.1
3-FullIns_4 405 3524 0.04 0.1 0.5 1.1 0.1 0.2 - 0.1
3-FullIns_5 2030 33751 0.02 5.4 14.8 23.2 5.4 32.2 - -
3-Insertions_3 56 110 0.07 1.0 7.0 0.2 0.1 0.1 - 2.3
3-Insertions_4 281 1046 0.03 - - - - - - -
3-Insertions_5 1406 9695 0.01 - - - - - - -
4-FullIns_3 114 541 0.08 0.1 0.1 0.3 0.1 0.1 0.1 0.1
4-FullIns_4 690 6650 0.03 0.1 1.9 2.3 0.1 0.8 - 0.1
4-FullIns_5 4146 77305 0.01 41.1 - 132.4 68.6 687.2 - -
4-Insertions_3 79 156 0.05 11.3 2300 0.3 0.3 0.2 - 2145
4-Insertions_4 475 1795 0.02 - - - - - - -
5-FullIns_3 154 792 0.07 0.1 0.1 0.3 0.1 0.1 0.1 0.1
5-FullIns_4 1085 11395 0.02 0.2 3.8 5.9 0.2 5.9 - 0.3
abb313GPIA 1557 53356 0.04 305.9 - 24.9 0.6 - - -
anna 138 493 0.05 0.1 0.1 0.2 0.1 0.1 0.1 0.1
ash331GPIA 662 4181 0.02 15.0 2.5 1.8 2.0 0.1 - 0.2
ash608GPIA 1216 7844 0.01 128.1 27.3 5.0 0.2 0.3 - -
ash958GPIA 1916 12506 0.01 1156 122.2 12.4 0.3 0.7 - 4.3
C2000.5 2000 999836 0.50 - - - - - - -
C2000.9 2000 1799532 0.90 - - - - - - -
C4000.5 4000 4000268 0.50 - - - - - - -
david 87 406 0.11 0.1 0.1 0.2 0.1 0.1 0.1 0.1
DSJC1000.1 1000 49629 0.10 - - - - - - -
DSJC1000.5 1000 249826 0.50 - - - - - - -
DSJC1000.9 1000 $449$ 0.90 - - - - - - -
DSJC125.1 125 736 0.09 2.4 0.8 0.3 1.7 0.1 1713 0.1
DSJC125.5 125 3891 0.50 - - - - - - -
DSJC125.9 125 6961 0.90 508.1 - - - - 7.0 -
DSJC250.1 250 3218 0.10 - - - - - - -
DSJC250.5 250 15668 0.50 - - - - - - -
DSJC250.9 250 27897 0.90 nan - - - - - -
DSJC500.1 500 12458 0.10 - - - - - - -
DSJC500.5 500 62624 0.50 - - - - - - -
DSJC500.9 500 112437 0.90 - - - - - - -
DSJR500.1 500 3555 0.03 0.1 0.1 1.5 0.1 0.1 270.3 0.1
DSJR500.1c 500 121275 0.97 3.1 - - 3.1 36.9 666.7 -
DSJR500.5 500 58862 0.47 79.0 - - - 31.0 3167 -
flat1000_50_0 1000 245000 0.49 - - - - - - -
flat1000_60_0 1000 245830 0.49 - - - - - - -
flat1000_76_0 1000 246708 0.49 - - - - - - -
flat300_20_0 300 21375 0.48 - - - - - 629.1 -
flat300_26_0 300 21633 0.48 - - - - - 1029 -
flat300_28_0 300 21695 0.48 - - - - - - -
fpsol2.i.1 496 11654 0.09 0.1 2.5 16.2 0.1 0.1 0.4 0.1
fpsol2.i.2 451 8691 0.09 0.1 1.0 11.7 0.1 0.1 0.4 0.1
fpsol2.i.3 425 8688 0.10 0.1 1.0 13.2 0.1 0.1 0.3 0.1
games120 120 638 0.09 0.1 0.1 0.3 0.1 0.1 0.1 0.1
homer 561 1629 0.01 0.1 0.1 0.9 0.1 0.1 0.1 0.1
huck 74 301 0.11 0.1 0.1 0.3 0.1 0.1 0.1 0.1
inithx.i.1 864 18707 0.05 0.1 3.5 21.4 0.1 0.2 1.2 0.1
inithx.i.2 645 13979 0.07 0.1 2.0 18.7 0.1 0.1 0.3 0.1
inithx.i.3 621 13969 0.07 0.1 1.9 19.3 0.1 0.1 0.3 0.1
jean 80 254 0.08 0.1 0.1 0.3 0.1 0.1 0.1 0.1
latin_square_10 900 307350 0.76 - - - - - - -
le450_15a 450 8168 0.08 21.8 13.2 4.1 10.1 0.2 - -
le450_15b 450 8169 0.08 10.6 28.1 4.0 0.2 0.2 - -
le450_15c 450 16680 0.17 696.1 - - - 58.7 - -
le450_15d 450 16750 0.17 289.8 - - - 45.0 - -
le450_25a 450 8260 0.08 0.1 0.5 4.9 0.1 0.1 1.4 0.1
le450_25b 450 8263 0.08 0.1 0.3 4.9 0.1 0.1 2.2 0.1
le450_25c 450 17343 0.17 - - - - - - -
le450_25d 450 17425 0.17 - - - - - - -
le450_5a 450 5714 0.06 6.9 30.3 2.0 0.1 0.1 - -
le450_5b 450 5734 0.06 5.3 34.5 1.9 0.1 0.1 - -
le450_5c 450 9803 0.10 2.9 5.6 3.0 0.1 0.1 - 0.1
le450_5d 450 9757 0.10 2.3 5.1 3.0 0.1 0.1 3330 11.2
miles1000 128 3216 0.40 0.1 0.5 5.8 0.1 0.1 0.1 0.1
miles1500 128 5198 0.64 0.1 1.9 21.9 0.1 0.1 0.1 0.1
miles250 128 387 0.05 0.1 0.1 0.3 0.1 0.1 0.1 0.1
miles500 128 1170 0.14 0.1 0.1 0.4 0.1 0.1 0.1 0.1
miles750 128 2113 0.26 0.1 0.2 2.3 0.1 0.1 0.1 0.1
mug100_1 100 166 0.03 0.6 0.1 0.3 0.6 0.1 0.6 -
mug100_25 100 166 0.03 0.6 0.1 0.3 0.5 0.1 0.5 -
mug88_1 88 146 0.04 0.4 0.1 0.3 0.3 0.1 0.3 371.7
mug88_25 88 146 0.04 0.4 0.1 0.3 0.4 0.1 0.4 241.7
mulsol.i.1 197 3925 0.20 0.1 0.8 5.0 0.1 0.1 0.1 0.1
mulsol.i.2 188 3885 0.22 0.1 0.4 5.4 0.1 0.1 0.1 0.1
mulsol.i.3 184 3916 0.23 0.1 0.4 5.6 0.1 0.1 0.1 0.1
mulsol.i.4 185 3946 0.23 0.1 0.4 5.7 0.1 0.1 0.1 0.1
mulsol.i.5 186 3973 0.23 0.1 0.4 5.7 0.1 0.1 0.1 0.1
myciel3 11 20 0.36 0.1 0.1 0.3 0.1 0.1 0.1 0.1
myciel4 23 71 0.28 0.1 0.1 0.3 0.1 0.1 5.4 0.1
myciel5 47 236 0.22 0.1 0.1 0.7 0.1 0.5 - 1.2
myciel6 95 755 0.17 0.1 0.1 1981 0.1 932.6 - -
myciel7 191 2360 0.13 0.1 0.5 - 0.1 - - -
qg.order100 10000 990000 0.02 - - - 1942 - - -
qg.order30 900 26100 0.06 138.0 0.2 21.9 0.3 3.5 - 0.6
qg.order40 1600 62400 0.05 1935 53.7 80.0 0.7 72.0 - -
qg.order60 3600 212400 0.03 1436 1710 2137 8.4 207.7 - -
queen10_10 100 1470 0.30 183.4 - 371.3 - - 170.3 -
queen11_11 121 1980 0.27 112.1 - 346.4 - - - -
queen12_12 144 2596 0.25 - - - - - - -
queen13_13 169 3328 0.23 - - - - - - -
queen14_14 196 4186 0.22 - - - - - - -
queen15_15 225 5180 0.21 - - - - - - -
queen16_16 256 6320 0.19 - - - - - - -
queen5_5 25 160 0.53 0.1 0.1 0.4 0.1 0.1 0.1 0.1
queen6_6 36 290 0.46 0.1 0.3 0.5 0.1 0.1 0.3 0.1
queen7_7 49 476 0.40 0.1 0.1 0.6 0.1 0.1 0.6 0.1
queen8_12 96 1368 0.30 0.3 0.1 1.0 0.1 0.1 7.7 0.1
queen8_8 64 728 0.36 1.3 5.8 3.4 0.4 12.7 5.2 1.0
queen9_9 81 1056 0.33 25.8 160.3 11.7 381.9 - 11.1 588.8
r1000.1 1000 14378 0.03 0.1 0.5 9.4 0.1 0.2 1.0 0.8
r1000.1c 1000 485090 0.97 - - - - - - -
r1000.5 1000 238267 0.48 - - - 847.5 1870 - -
r125.1 125 209 0.03 0.1 0.1 0.3 0.1 0.1 0.1 0.1
r125.1c 125 7501 0.97 0.1 3.5 37.9 0.1 0.1 0.1 0.1
r125.5 125 $38$ 0.50 4.3 2.7 8.9 0.3 0.1 11.6 0.1
r250.1 250 867 0.03 0.1 0.1 0.5 0.1 0.1 0.1 0.1
r250.1c 250 30227 0.97 0.1 40.3 103.3 0.1 0.3 30.7 1.0
r250.5 250 14849 0.48 94.5 36.0 71.1 3.0 2.6 185.9 199.6
school1 385 19095 0.26 0.3 3.1 18.2 0.3 0.6 968.9 5.7
school1_nsh 352 14612 0.24 0.1 4.8 11.6 0.1 0.1 1840 6.0
wap01a 2368 110871 0.04 - - 1778 - 124.1 - -
wap02a 2464 111742 0.04 - - 1054 - 144.4 - -
wap03a 4730 286722 0.03 - - - - - - -
wap04a 5231 294902 0.02 - - - - - - -
wap05a 905 43081 0.11 0.1 5.6 61.2 0.1 0.7 5.2 0.9
wap06a 947 43571 0.10 413.0 - 114.0 575.5 31.6 - -
wap07a 1809 103368 0.06 - - - - - - -
wap08a 1870 104176 0.06 - - 504.9 2345 - - -
will199GPIA 701 6772 0.03 0.3 1.2 3.2 0.9 0.2 1.5 0.3
zeroin.i.1 211 4100 0.19 0.1 0.9 5.5 0.1 0.1 0.1 0.1
zeroin.i.2 211 3541 0.16 0.1 0.4 4.3 0.1 0.1 0.1 0.1
zeroin.i.3 206 3540 0.17 0.1 0.4 4.5 0.1 0.1 0.1 0.1